cos^3a sin3a+sin^3a cos3a=3/4sin4a
1
Expert's answer
2019-05-02T02:36:57-0400
"\text{sin}^3\alpha \text{cos}(3\alpha) + \text{cos}^3\alpha \text{sin}(3\alpha)="

"=\text{sin}^3\alpha(4\text{cos}^3\alpha-3\text{cos}\alpha)+\text{cos}^3\alpha(3\text{sin}\alpha-4\text{sin}^3\alpha)="

"=4\text{sin}^3\alpha\text{cos}^3\alpha-3\text{sin}^3\alpha\text{cos}\alpha+3\text{cos}^3\alpha\text{sin}\alpha-4\text{cos}^3\alpha\text{sin}^3\alpha="

"=3\text{sin}\alpha\text{cos}\alpha(\text{cos}^2\alpha-\text{sin}^2\alpha)="

"=\frac{3}{2}\text{sin}(2\alpha)\text{cos}(2\alpha)=\frac{3}{4}\text{sin}(4\alpha)"